Draw the expected structures for compounds with the following empirical formulas that would fulfill the 18-electon guideline for electronic saturation..               a)      Cp2Os(CO)2    ...

  1.        Draw the expected structures for compounds with the following empirical formulas that would fulfill the 18-electon guideline for electronic saturation..

              a)      Cp2Os(CO)2

              b)    η 5-Cp W(CO)2
Answer #1

First I want to tell you that these compound don't follow 18 Electrons rule.

Electrons contribution of CO = 2

Electrons contribution of CP (cyclopentadiene) = 5

Valence electrons of Os(osmium) = 8

Valence electrons of w(tungsten) = 6

a) CP2Os(CO)2

(5×2) + 8 + (2×2)

= 10+8+4 = 22 Electrons.

Hence, this compound don't follow 18 elecElect rule.

Because central metal atom is 1 so total electrons count should be 18, rather than 22.

b) n5-CP W(CO)2

5 + 6 + (2×2)

= 5 +6+4 = 15

Hence, this compound don't follow 18 Electrons rule.

Because here central metal atom is 1 so total Electrons should be 18 rather than 15.

Suppose if here are 2 central metal atoms then total Electrons count should be 18×2=36, according to 18 Electrons rule.

Expected structure of these shown below-

2. coo CO Co Co OS nelpatici

Hepticity :- it represents how many no. of carbon atoms of a cyclic ring attached to central metal atom.

In second compound hepticity is 5 that means 5 carbon atoms attached to central metal atom.